# Broadcast in Almost Mixing Time

### v2

@@ 3.2 Coalescing-Branching Random Walk
We now overview the main primitive for our algorithm for an Erdős–Rényi graph – the COalescing-BRAnching Random Walk (COBRA walk). COBRA walk was first introduced by Dutta et al. [15] in their work “Coalescing-Branching Random Walks on Graph” [15], with subsequent refinements presented in [12, 41, 6]. The COBRA walk is a generalization of the classical random walk, defined as follows: At round 00, a source node $s\in V$ possesses a token. At round $r$, each node possessing a token selects $\kappa$ of its neighbors uniformly at random, sends a token copy to each of them, and these neighbors are said to possess a token at round $r+1$. Here, $\kappa$, referred to as the branching factor, can be generalized to any positive real number [12]. When $\kappa=1$, the COBRA walk reduces to the classical random walk. From now on, we consider $\kappa$ to always be $2$. It is important to note that if a node receives multiple token copies in a round, it behaves as if it has received only one token; it will still choose $\kappa$ neighbors uniformly at random. This property, where received token copies coalesce at a node, gives the primitive its name.
Cooper et al. [12] studied the cover time $T$ of the COBRA walk on regular expanders and obtained the following theorem which we use in our result
@@
Let $G$ be a connected $n$-vertex regular graph. Let $\lambda_{2}$ be the second largest eigenvalue (in the absolute value) of the normalized adjacency matrix of $G$. Then the following holds with probability at least $1-O(\frac{1}{n^{2}})$.
$T=O\left(\frac{\log n}{(1-\lambda_{2})^{3}}\right)$
@@ 3.3 Random Graph Algorithm Description
    Constructing Spanning Subgraphs through Multiple COBRA Walks: The source node $s$ initiates $\delta$ COBRA walks by creating $\delta$ tokens labeled $1$ through $\delta$. When a token from the $i$-th COBRA walk is sent along an edge $e:\ (u,v)$, nodes $u$ and $v$ mark $e$ as part of the $i$-th subgraph. Note that a single edge may belong to multiple subgraphs.
    When running multiple COBRA walks simultaneously, congestion can occur if a node attempts to send multiple tokens from different COBRA walks along the same edge in a single round. Since only one token can traverse an edge per round, this creates a bottleneck that needs to be managed. To address this, we organize the process into phases, where each phase consists of $2$ rounds. In each phase, spanning rounds $\{2r,2r+1\}$, every node $u$ distributes two tokens for each COBRA walk whose token(s) it received during the previous phase. Since there are $\delta$ COBRA walks, node $u$ could have received at most $p\leq\delta$ distinct tokens. Let these tokens be denoted by $t_{1},\ldots,t_{p}$. Node $u$ then distributes each token twice, as follows: it enumerates its neighbors as $v_{1},\ldots,v_{\ell}$, with $\ell\geq\delta$, generates two independent random permutations $\sigma_{1},\sigma_{2}\in S_{\ell}$, and sends token $t_{i}$ to neighbor $v_{\sigma_{1}(i)}$ in round $2r$ and to neighbor $v_{\sigma_{2}(i)}$ in round $12r+$ 1. Consequently, for any fixed COBRA walk $i$ and any node $u$, the token belonging to the $i$-th COBRA walk is sent to a neighbor of $u$ chosen uniformly at random, independently of the other tokens in that same COBRA walk. (Different cobra walks, however, need not be independent.)
@@
    Constructing Tree Packings: The source $s$ initiates a BFS on each subgraph to transform it into a tree. By the end of this step, the algorithm constructs a tree packing $\{T_{i}\}_{i\in[\delta]}$. This step takes the number of phases that is at most the height of the highest tree, that is at most $O(T)$ phases, and hence $2T\cdot O(T)=O(T^{2})$ rounds.
@@
    Distributing Messages: The source node $s$ evenly divides the set of messages $M$ across the $\delta$ trees, ensuring that each tree receives $\frac{k}{\delta}$ messages. These messages are then downcasted along the trees one by one. To broadcast $\frac{k}{\delta}$ messages in each tree of height $O(T)$ in parallel, one needs $O(\frac{k}{\delta}+T)$ phases, that is $O(T\cdot(\frac{k}{\delta}+T))$ rounds.
@@ 4.1 Proof Outline
The proof proceeds in three main steps. First, we argue that making a random graph regular by adding self-loops does not significantly affect its expansion properties. To this end, we rely on a result of Hoffman et al. [33], which shows that a random graph is a good expander with high probability. We also use standard Chernoff bound arguments to establish that random graphs are nearly regular, meaning the ratio between maximum and minimum degrees is close to one. Finally, we invoke a classical consequence of Weyl’s inequality, which ensures that small perturbations to the diagonal of a matrix only slightly affect its eigenvalues. Together, these ingredients imply that regularizing the graph preserves its spectral expansion up to a small error.
@@
The second step is to prove that each individual COBRA walk successfully covers the entire network. The permutation trick allows us to claim that though COBRA walks are not independent, their marginal distributions stay as if they were. Using this, Theorem 3.7 by Cooper et al., together with a union bound, guarantees that all COBRA walks complete successfully within $O(\log n)$ phases w.h.p.
@@
In the final step, we argue that the algorithm produces a tree packing with size $\delta(G)$, height $O(\log n)$, and weight $O(\log n)$. This tree packing allows for broadcasting all messages in $O(\log^{2}n+\log n\cdot\frac{k}{\delta(G)})$ rounds. This is optimal up to additive $\log^{2}n$ and multiplicative $\log n$, as $\frac{k}{\delta(G)}$ provides a natural lower bound for the broadcast time.
@@ 4.2 Introducing Regularity while Maintaining Expansion
We start by providing relevant concepts from spectral theory.
@@
Proof 4.5.
Let us fix a vertex $v$ and consider the number of its incident edges. For each potential edge $e_{i}$, $i\in[n-1]$ let us introduce an indicator variable $\chi_{i}$ which is equal to $1$ if the edge exists and to 00 if it does not. The number of edges $v$ has is then $\sum\limits_{i\in[n-1]}\chi_{i}$. The expectation of that is $p(n-1)$, and applying Chernoff bounds, we get the following bounds on the degree of $v$
@@
The next ingredient is to show that the slight perturbation of diagonal elements of the matrix induces only a little change on its eigenvalues.
@@
Let $A$ be an $n\times n$ adjacency matrix of a connected graph and let $D$ be a diagonal matrix such that $D_{ii}=\sum\limits_{j\in[n]}A_{ij}$. Let $E$ be a $n\times n$ diagonal matrix such that $0\leq E_{ii}\leq\varepsilon$ for some $10<\varepsilon<$ 1 and all $i\in[n]$.
@@
Proof 4.7 (Proof sketch.).
The idea of the proof is to express $\overline{A+DE}$ as a sum of $\overline{A}$ and matrices with the small spectral norms, and then apply a corollary of Weyl’s Theorem [54], that is for $n\times n$ matrices $M_{1}$ and $M_{2}$, it holds that
$|\lambda_{2}(M_{1}+M_{2})-\lambda_{2}(M_{1})|\leq||M_{2}||_{2}$
@@
Finally, using Lemmas 4.4 and 4.6 alongside Theorem 4.3, we show that w.h.p., regularizing an Erdős–Rényi graph $G(n,p)$ by adding self-loops for every node to reach $\Delta(G)$ does not ruin its expansion properties.
Lemma 4.8.
Proof 4.9.
@@
4.3 Success of Multiple COBRAs
Let us recall the permutation trick used to avoid congestion among multiple COBRA walks. To be able to run multiple COBRA walks in parallel, we partition the execution into phases, each lasting $2$ rounds. In each phase, spanning rounds ${2r,2r+1}$, every node $u$ distributes two tokens for each COBRA walk whose token(s) it received during the previous phase. Since there are $\delta$ COBRA walks, node $u$ could have received at most $p\leq\delta$ distinct tokens. Let these tokens be denoted by $t_{1},\ldots,t_{p}$. Node $u$ then distributes each token twice, as follows: it enumerates its neighbors as $v_{1},\ldots,v_{\ell}$, with $\ell\geq\delta$, generates two independent random permutations $\sigma_{1},\sigma_{2}\in S_{\ell}$, and sends token $t_{i}$ to neighbor $v_{\sigma_{1}(i)}$ in round $2r$ and to neighbor $v_{\sigma_{2}(i)}$ in round $12r+$ 1.
Claim 1.
To conclude the analysis of multi-COBRA’s performance on a random graph, we compile our knowledge from sections 4.2 and 4.3 and get the following lemma.
Lemma 4.10.
If the initial network graph is an Erdős–Rényi graph $G(n,p)$ with $p\geq\frac{C_{p}\log n}{n}$, all COBRA walk cover the graph in $O(T)$ rounds with probability at least $1-O(\frac{1}{n})$.
Proof 4.11.
By Lemma 4.8 we know that $G^{\prime}$ the graph we obtain from $G$ after adding self-loops - has $1-\lambda_{2}(G^{\prime})\geq\frac{1}{14}$ with probability at least $1-O(\frac{1}{n^{2}})$. Therefore, according to Theorem 3.7, a COBRA walk succeeds to cover $G^{\prime}$ in $O\left(\frac{\log n}{(1-\lambda_{2}(G^{\prime}))^{3}}\right)=O(T)=O(\log n)$ rounds with probability at least $1-O(\frac{1}{n^{2}})$. And Claim 1 tells us that while the COBRA walks are not independent, the marginal distribution of each individual process matches the distribution it would have under independent execution. Hence, we can apply a union bound and conclude that $\delta(G)\leq n$ COBRA walks cover $G^{\prime}$ in $O(\log n)$ phases with probability at least $1-O(\frac{1}{n})$.
4.4 Tree Packing and Broadcast
Lemma 4.12.
Proof 4.13.
Lemma 4.14.
Proof 4.15.
In the rest of this section, we will run protocols on all the subgraphs in parallel. To achieve this despite potential congestion (recall that each edge may belong to multiple subgraphs), we again organize the execution into phases. Now, each phase spans $O(\log n)$ rounds, ensuring that messages sent along any shared edge are distributed across the protocols without conflict. Specifically, if a protocol would send a message along an edge in a particular round when executed independently, all such messages from different subgraphs are scheduled within the same phase. This phased execution allows all protocols to proceed in parallel while respecting the edge capacity. As a result, the combined round complexity of the protocols increases by at most a factor of $O(\log n)$ compared to running an individual protocol.
We are now ready to prove Theorem 1.4. See 1.4
Proof 4.16 (Proof of Theorem 1.4).
See 1.3
Proof 4.17 (Proof of Theorem 1.3).
First, build a tree packing from Theorem 1.4. Then, $s$ evenly distributes messages among the obtained trees, so that each tree receives $\frac{k}{\delta}$ messages. After that, in each tree, nodes downcast corresponding messages. The algorithm for downcasting messages $\{m_{1},\ldots,m_{k}\}$ from the root of a single tree works as follows. In the first round, the root sends the first message ($m_{1}$) to all its immediate children. In the second round, the children forward $m_{1}$ to their respective children (the root’s grandchildren), while the root simultaneously sends the second message ($m_{2}$) to its immediate children. This process continues iteratively: in each subsequent round, the root sends the next message ($m_{i}$) to its children, and all other nodes forward the message they received in the previous round to their respective children. This way, for $k^{\prime}$ messages and a tree of height $h$ it takes $h+k^{\prime}-1$ rounds for every node to discover every message.
Multiplying by a congestion factor of $O(\log n)$, we get that the round complexity of broadcasting $k$ messages in $\delta(G)$ spanning trees of height $O(\log n)$ is
$O\left(\log^{2}n+\log n\cdot\frac{k}{\delta(G)}\right).$
Remark 4.18.
@@ 5.1 NP-Hardness
In our reduction, for simplicity of presentation, we allow edges to have arbitrary bandwidth instead of $\tilde{O}(1)$, since, as we show, this can be simulated in CONGEST. In the reduction, the initial set $S$ corresponds to the set $M$ of messages, and $s$ has two dedicated children to which it can send $n_{1}$ and $n_{2}$ messages respectively with $n_{1}+n_{2}=k$, simulating the splitting. Deciding how to split messages between these two children is the only “smart” choice an algorithm should make; all other nodes are just forwarding messages they receive. For the full version of the proof, please see Appendix C.

### v3

@@ 3.2 Coalescing-Branching Random Walk
We now overview the main primitive for our algorithm for an Erdős–Rényi graph – the COalescing-BRAnching Random Walk (COBRA walk). COBRA walk was first introduced by Dutta et al. [13] in their work “Coalescing-Branching Random Walks on Graph” [13], with subsequent refinements presented in [12, 38, 5]. The COBRA walk is a generalization of the classical random walk, defined as follows: At round 0, a source node $s\in V$ possesses a token. At round $r$, each node possessing a token selects $\kappa$ of its neighbors uniformly at random, sends a token copy to each of them, and these neighbors are said to possess a token at round $r+1$. Here, $\kappa$, referred to as the branching factor, can be generalized to any positive real number [12]. When $\kappa=1$, the COBRA walk reduces to the classical random walk. From now on, we consider $\kappa$ to always be $2$. It is important to note that if a node receives multiple token copies in a round, it behaves as if it has received only one token; it will still choose $\kappa$ neighbors uniformly at random. This property, where received token copies coalesce at a node, gives the primitive its name.
Cooper et al. [12] studied the cover time of the COBRA walk on regular expanders and obtained the following theorem, which we use in our result
@@
Let $G$ be a connected $n$-vertex regular graph. Let $\lambda_{2}$ be the second largest eigenvalue (in the absolute value) of the normalized adjacency matrix of $G$. Then after $O\left(\frac{\log n}{(1-\lambda_{2})^{3}}\right)$ steps the COBRA walk covers $G$ with probability $1-O(\frac{1}{n^{2}})$.
@@ 3.3 Random Graph Algorithm Description
    Constructing Spanning Subgraphs through Multiple COBRA Walks: The source node $s$ initiates $\delta$ COBRA walks by creating $\delta$ tokens labeled $1$ through $\delta$. When a token from the $i$-th COBRA walk is sent along an edge e:(u,v)e:\ (u,v), nodes $u$ and $v$ mark $e$ as part of the $i$-th subgraph. Note that a single edge may belong to multiple subgraphs.
    When running multiple COBRA walks simultaneously, congestion can occur if a node attempts to send multiple tokens from different COBRA walks along the same edge in a single round. Since only one token can traverse an edge per round, this creates a bottleneck that needs to be managed. To address this, we organize the process into phases, where each phase consists of $2$ rounds. In each phase, spanning rounds $\{2r,2r+1\}$, every node $u$ distributes two tokens for each COBRA walk whose token(s) it received during the previous phase. Since there are $\delta$ COBRA walks, node $u$ could have received at most $p\leq\delta$ distinct tokens. Let these tokens be denoted by $t_{1},\ldots,t_{p}$. Node $u$ then distributes each token twice, as follows: it enumerates its neighbors as $v_{1},\ldots,v_{\ell}$, with $\ell\geq\delta$, generates two independent random permutations $\sigma_{1},\sigma_{2}\in S_{\ell}$, and sends token $t_{i}$ to neighbor $v_{\sigma_{1}(i)}$ in round $2r$ and to neighbor $v_{\sigma_{2}(i)}$ in round $2r+1$. Consequently, for any fixed COBRA walk $i$ and any node $u$, the token belonging to the $i$-th COBRA walk is sent to a neighbor of $u$ chosen uniformly at random, independently of the other tokens in that same COBRA walk. (Different cobra walks, however, need not be independent.)
@@
    Constructing Tree Packings: The source $s$ initiates a BFS on each subgraph to transform it into a tree. By the end of this step, the algorithm constructs a tree packing $\{T_{i}\}_{i\in[\delta]}$. This step takes the number of phases that is at most the maximal eccentricity of $s$ among all the subgraphs, that is at most $O(T)$ phases, and hence $2T\cdot O(T)=O(T^{2})$ rounds.
@@
    Distributing Messages: The source node $s$ evenly divides the set of messages $M$ across the $\delta$ trees, ensuring that each tree receives $\frac{k}{\delta}$ messages. These messages are then downcasted along the trees one by one. To broadcast $\frac{k}{\delta}$ messages in a single tree of diameter $O(T)$ one needs $O(\frac{k}{\delta}+T)$ rounds. Hence, doing it in parallel in all trees takes $O(T\cdot(\frac{k}{\delta}+T))$ rounds.
@@ 4.1 Proof Outline
The proof proceeds in three main steps. First, we argue that making a random graph regular by adding self-loops does not significantly affect its expansion properties. To this end, we rely on a result of Hoffman et al. [32], which shows that a random graph is a good expander with high probability. We also use standard Chernoff bound arguments to establish that random graphs are nearly regular, meaning the ratio between maximum and minimum degrees is close to one. Finally, we invoke a classical consequence of Weyl’s inequality, which ensures that small perturbations to the diagonal of a matrix only slightly affect its eigenvalues. Together, these ingredients imply that regularizing the graph preserves its spectral expansion up to a small error. We discuss the details at the end of this section, namely in subsection 4.4.
@@
Assuming the expansion properties remain, the second step is to prove that each individual COBRA walk successfully covers the entire network. The permutation trick allows us to claim that, though COBRA walks are not independent, their marginal distributions stay as if they were. Using this, Theorem 3.7 by Cooper et al., together with a union bound, guarantees that all COBRA walks cover the whole graph within $O(\log n)$ phases w.h.p.
@@
In the final step, we argue that the algorithm produces a tree packing with size $\delta(G)$, diameter $O(\log n)$, and weight $O(\log n)$. This tree packing allows for broadcasting all messages in $O(\log^{2}n+\log n\cdot\frac{k}{\delta(G)})$ rounds. This is optimal up to additive $\log^{2}n$ and multiplicative $\log n$, as $\frac{k}{\delta(G)}$ provides a natural lower bound for the optimal broadcast time: if there is a node with degree $\delta$, it needs at least $k/\delta$ rounds to receive $k$ messages.
We now give the detailed proof starting from step 2, assuming step 1.
4.2 Success of Multiple COBRAs
In this section, we demonstrate that multiple COBRA walks cover the graph fast, provided that it retains its expansion properties after adding self-loops. Formally, we assume the following lemma, which we prove in section 4.4.
Lemma 4.1.
Now, to see why multiple COBRAs do not congest, let us recall the permutation trick used in the algorithm. To be able to run multiple COBRA walks in parallel, we partition the execution into phases, each lasting $2$ rounds. In each phase, spanning rounds 2r,2r+1{2r,2r+1}, every node $u$ distributes two tokens for each COBRA walk whose token(s) it received during the previous phase. Since there are $\delta$ COBRA walks, node $u$ could have received at most $p\leq\delta$ distinct tokens. Let these tokens be denoted by $t_{1},\ldots,t_{p}$. Node $u$ then distributes each token twice, as follows: it enumerates its neighbors as $v_{1},\ldots,v_{\ell}$, with $\ell\geq\delta$, generates two independent random permutations $\sigma_{1},\sigma_{2}\in S_{\ell}$, and sends token $t_{i}$ to neighbor $v_{\sigma_{1}(i)}$ in round $2r$ and to neighbor $v_{\sigma_{2}(i)}$ in round $2r+1$.
Claim 1.
To conclude the analysis of multi-COBRA’s performance on a random graph, we combine Lemma 4.1 and Claim 1 and get the following lemma.
Lemma 4.2.
If the initial network graph is an Erdős–Rényi graph $G(n,p)$ with $p\geq\frac{C_{p}\log n}{n}$, all COBRA walks cover the graph in $O(T)$ rounds with probability at least $1-O(\frac{1}{n})$.
Proof 4.3.
By Lemma 4.1 we know that $G^{\prime}$ - the graph we obtain from $G$ after adding self-loops - has $1-\lambda_{2}(G^{\prime})\geq\frac{1}{14}$ with probability at least $1-O(\frac{1}{n^{2}})$. Therefore, according to Theorem 3.7, a COBRA walk succeeds to cover $G^{\prime}$ in $O\left(\frac{\log n}{(1-\lambda_{2}(G^{\prime}))^{3}}\right)=O(T)=O(\log n)$ rounds with probability at least $1-O(\frac{1}{n^{2}})$. And Claim 1 tells us that while the COBRA walks are not independent, the marginal distribution of each individual process matches the distribution it would have under independent execution. Hence, we can apply a union bound and conclude that $\delta(G)\leq n$ COBRA walks cover $G^{\prime}$ in $O(\log n)$ phases with probability at least $1-O(\frac{1}{n})$.
4.3 Tree Packing and Broadcast
Lemma 4.4.
Proof 4.5.
Lemma 4.6.
Proof 4.7.
In the rest of this section, we will analyze protocols that run on all the subgraphs in parallel. To achieve this parallelism despite potential congestion (recall that each edge may belong to multiple subgraphs), the execution is again organized into phases. Now, each phase spans $O(\log n)$ rounds, ensuring that messages sent along any shared edge are distributed across the protocols without conflict. Specifically, messages that would be sent on an edge in the same round by independently executed protocols are all scheduled within a single phase across subgraphs. This phased execution allows all protocols to proceed in parallel while respecting the edge capacity. As a result, the combined round complexity of the protocols increases by at most a factor of $O(\log n)$ compared to running an individual protocol.
We are now ready to prove Theorem 1.4. See 1.4
Proof 4.8 (Proof of Theorem 1.4).
See 1.3
Proof 4.9 (Proof of Theorem 1.3).
First, build a tree packing from Theorem 1.4. Then, $s$ evenly distributes messages among the obtained trees, so that each tree receives $\frac{k}{\delta}$ messages. After that, in each tree, nodes downcast corresponding messages. The algorithm for downcasting messages $\{m_{1},\ldots,m_{k}\}$ from the root of a single tree works as follows. In the first round, the root sends the first message ($m_{1}$) to all its immediate children. In the second round, the children forward $m_{1}$ to their respective children (the root’s grandchildren), while the root simultaneously sends the second message ($m_{2}$) to its immediate children. This process continues iteratively: in each subsequent round, the root sends the next message ($m_{i}$) to its children, and all other nodes forward the message they received in the previous round to their respective children. This way, for $k^{\prime}$ messages and a tree of diameter $H$ it takes $H+k^{\prime}-1$ rounds for every node to discover every message.
Multiplying by a congestion factor of $O(\log n)$, we get that the round complexity of broadcasting $k$ messages in $\delta(G)$ spanning trees of diameter $O(\log n)$ is
$O\left(\log^{2}n+\log n\cdot\frac{k}{\delta(G)}\right).$
Remark 4.10.
@@ 4.4 Introducing Regularity while Maintaining Expansion
In this Section, we prove Lemma 4.1 that states that graph’s expansion properties are preserved after adding self-loops. We start by providing relevant concepts from spectral theory.
@@
Proof 4.15.
Let us fix a vertex $v$ and consider the number of its incident edges. For each potential edge $e_{i}$, $i\in[n-1]$ let us introduce an indicator variable $\chi_{i}$ which is equal to $1$ if the edge exists and to 0 if it does not. The number of edges $v$ has is then $\sum\limits_{i\in[n-1]}\chi_{i}$. The expectation of that is p(n1)p(n-1), and applying Chernoff bounds, we get the following bounds on the degree of $v$
@@
The next ingredient is to show that the slight perturbation of the diagonal elements of the matrix induces only a little change in its eigenvalues.
@@
Let $A$ be an $n\times n$ adjacency matrix of a connected graph and let $D$ be a diagonal matrix such that $D_{ii}=\sum\limits_{j\in[n]}A_{ij}$. Let $E$ be a $n\times n$ diagonal matrix such that $0\leq E_{ii}\leq\varepsilon$ for some $0<\varepsilon<1$ and all $i\in[n]$.
@@
Proof 4.17 (Proof sketch.).
The idea of the proof is to express $\overline{A+DE}$ as a sum of $\overline{A}$ and matrices with the small spectral norms, and then apply a corollary of Weyl’s Theorem [48], that is, for $n\times n$ matrices $M_{1}$ and $M_{2}$, it holds that
[…]
@@
Finally, using Lemmas 4.14 and 4.16 alongside Theorem 4.13, we show that w.h.p., regularizing an Erdős–Rényi graph $G(n,p)$ by adding self-loops for every node to reach $\Delta(G)$ does not ruin its expansion properties.
See 4.1
Proof 4.18.
@@ 5.1 NP-Hardness
In our reduction, for simplicity of presentation, we allow edges to have arbitrary bandwidth instead of $\tilde{O}(1)$, since, as we show, this can be simulated in CONGEST. In the reduction, the initial set $S$ corresponds to the set $M$ of messages, and $s$ has two dedicated children to which it can send $n_{1}$ and $n_{2}$ messages, respectively with $n_{1}+n_{2}=k$, simulating the splitting. Deciding how to split messages between these two children is the only “smart” choice an algorithm should make; all other nodes are just forwarding messages they receive. For the full version of the proof, please see Appendix C.